# HARNESSING Nth ROOT GATES FOR ENERGY STORAGE

### v1

@@ IV.2 Comparison between circuits with Full CNOT and NRCG
NCRGs can be interpreted as a stepped application of a full CNOT gate as explained more thoroughly in section II.2. We then wish to compare if applying fractions of CNOTs instead of a full CNOT is advantageous in the present case. We compare the ergotropy over 2 cycles in which iterations have either of one CNOT or N = 15 steps of NRCG, magenta. For each panel we consider $\theta$ such that the ergotropy of the system considered in the panel reaches its maximum when N = 15. The top left and top center panels in Fig. 3, showing case 1 and 2, for the two qubit systems operate with only a single gate per iteration. These particular systems demonstrate well how the NRCGs operate as a slower application of the full CNOT gate with the systems’ ergotropy reaching the same values for both types of gates after 1 cycle. Circuits with the full CNOT gate allows to reach only two specific values of ergotropy. However, NRCGs allow access to these and to all intermediate values. Such fine-tuning (increase and decrease) of ergotropy with the system evolution could prove advantageous when a particular value of ergotropy or the corresponding intermediate state are desired, that are in-between the initial and final ergotropy and states generated by a full CNOT gate.
@@
With the exception of cases 1 and 2 for two-qubit systems, NRCG-based cycles are not equivalent to cycles from corresponding circuits with CNOTs even when considering the end results, see Fig. 3. Apart from easy-reach of intermediate values of ergotropy, we can see other two ways in which the use of NRCGs could be advantageous: (i) at points where the ergotropy of the total system has a larger value than that of the standard CNOT gate within the same cycle; (ii) when the maximum ergotropy reached using NRCGs is higher that the one of the corresponding CNOT-gate circuits. The first possibility can be observed in all panels of Fig. 3; the second can be seen in the three-qubit cases 1 and 2.
@@
Similar results are found for the case of qubit B inizialized in $\Phi=0$ (not shown). We note that a larger value of the ergotropy may be reached after the two cycles considered in Fig. 3 (not shown).
@@ IV.3 Ergotropy variation for two- and three-qubit systems
Our systems do not start from a passive state, so the ergotropy $W_{max}$ at the end of the process may be influenced by the initial one. This is not a simple relation, as the systems studied undergo ’discharging’ and ’charging’ processes as the number of iterations increase, as discussed below.
In addition, when considering $W_{max}$ (Fig. 2), a direct comparison between the two and three-qubit systems is not easy to perform: when all common parameters are the same, the three-qubit system inherently start from more energy and ergotropy than the two-qubit system by merit of having more qubits. For the dynamics considered, this almost always translates into acquiring more ergotropy by the end of the process (compare first and second row in Fig. 2). From this we can argue that looking at the maximum ergotropy may not be a fair comparison between systems with different quantities of qubits. For these reasons, we introduce the ergotropy variation ($\Delta W_{max}$) which is defined by the following equation
@@
where $W_{max}^{(f)}$ is the ergotropy of the final state and $W_{max}^{(0)}$ is the ergotropy of the initial state of the total system. Equation (9) allows the identification of two regimes with respect to the initial state, charging or discharging. A charging regime is characterised by a positive value of ergotropy variation and a discharging regime is characterised by a negative value of ergotropy variation. $\Delta W_{max}$ allows for a clearer comparison across different systems. For example an equal or greater positive change in the two-qubit system compared to the corresponding three-qubit system may suggest that an easier-to-implement two-qubit circuit may be more suited to accrue energy as a battery.
Figure 4: Ergotropy variation $\Delta W_{max}$ for $\phi=0$. Else, same parameters as in figure 2.
Figure 5: Ergotropy variation $\Delta W_{max}$ for $\phi=\pi$. Same parameters as in figure 2.
$\Delta W_{max}$ is evaluated and plotted for $\phi=0$, see Fig. 4 and $\phi=\pi$, see Fig. 5. The corresponding panels for three-qubit systems show a dramatically different landscape, with the maxima no longer being constrained to $\theta\approx\pi$. We can see for case 2 that for the two and three-qubit system $\Delta W_{max}$ is small when compared to case 1 and 3 showing that it is not the optimal case for increasing or decreasing ergotropy from the initial state of the system. Cases 1 and 3 in Fig. 4 display much larger peaks and troughs for $\Delta W_{max}$ with respect to case 2, indicating a stronger performance as a quantum battery with well defined charging and discharging regimes. Comparison of cases 1 and 2, which correspond to inverted gates, indicates that having qubit A (thermalized initially at the higher temperature) being a control qubit produces more favourable changes in ergotropy. Also, the introduction of an extra 2 gates per iteration in case 3 fails to provide higher maxima in $\Delta W_{max}$ than case 1. Across all panels in Fig. 4 the charging regions (positive $\Delta W_{max}$) are dominating over discharging region, which is of advantage when looking to optimal conditions for operating a quantum battery. When qubit B is initialised with $\phi=\pi$ for three-qubit systems and cases 1 and 3 the maximum variation of ergotropy changes from 0.49 $\epsilon_{2B}$ to 0.62 $\epsilon_{2B}$, which is a 26.5% improvement in performance. For most cases, the maximum $\Delta W_{max}$ regions now lie closer to $\theta=\pi/2$. This shows that the extra initial coherence enhances the charging performance in most circuits considered, and especially so for case 1 and 3 and three-qubit systems.
@@ IV.4 Ergotropy Ratio for two- and three-qubit circuits
Figure 6: Ergotropy ratio $W_{ratio}$ for $\Phi=\pi$. Same parameters as in figure 2.
Here we introduce the ergotropy ratio which is defined as,
@@
where $W_{max}$ is the ergotropy (eq.7) and $Tr[\rho^{(f)}H]$ is the internal energy of the system in its final state. The ergotropy ratio takes a value between 0 and 1 where a value closer to 1 signifies a larger quantity of extractable energy. The ergotropy ratio is good measure of how efficiently a battery is performing[36]. With this in mind, here $W_{ratio}$ is introduced to look for optimal initial conditions for an efficient quantum battery.
Fig. 6 shows the ergotropy ratio for the total system for the same parameters as figure 2, in particular qubit B is initialised with $\Phi=\pi$. For all cases we obtain a greater value of ergotropy ratio when qubit B is initialised with a large value of $\theta$, but not necessarily $\theta=\pi$. The maximum value for $W_{ratio}$ across all panels in Fig. 6 are similar, suggesting no particular advantage with using a three instead of a two-qubit system. When initialising qubit B with $\Phi=\pi$ (see Fig. 6) we observe a marked increased in areas of high $W_{ratio}$ for values of $\theta$ closer to $\pi/2$ for both two (case 3) and three (cases 1 and 3) qubit systems, supporting advantage from a larger initial coherence. For comparison, results with qubit B initialized at $\Phi=0$ are shown in Fig. 14, appendix C.
While $W_{ratio}$ is a good indicator for which initial conditions lead to a larger percentage of total energy that can be extracted as work, it does not show if that large percentage also corresponds to a region of high charging: indeed $W_{ratio}$ and $\Delta W_{max}$ have quite different behaviours (e.g. compare Figs. 4 and 14). Ideally, one would like to optimize both quantities at the same time to quantitatively assess how well the initial conditions, system, and circuit considered perform as a battery. In the next section, we will introduce a related figure of merit.
@@ IV.5 Figure of Merit for optimal battery, two- and three-qubit systems
Figure 7: Figure of merit, FoM, with $\Phi=\pi$. Same parameters as figure 2
The figure of merit we propose to find the optimal circuit, number of iterations and initial conditions for charging a quantum battery has the form
@@
This combination of quantities shows where the best balance of energy that can be extracted (ergotropy ratio) and ergotropy increase from the initial condition (ergotropy difference) lies, giving a clearer indication of the optimal system.
In Fig. 7 we plot $FoM$ for the same parameter scans of Figs. 5 and 6. In particular, qubit B is initialized with $\Phi=\pi$. The results corresponding to $\Phi=0$ are plotted in Appendix C, Fig. 15. The topography of these graphs tends to mirror that of the variation in ergotropy, in general more variable with iterations than that the ergotropy ratio. Interestingly, when $\Phi=0$ (Fig. 15) we see the highest FoM value regions when $\theta\approx 0$ and $\theta\approx\pi$, suggesting that the system performs better as a battery when qubit B is initialised closer to a state with no coherences ($\Phi=0$, with $\theta=0$ or $\pi$). When comparing this to Fig. 7 ($\Phi=\pi$) the opposite is true, with the greatest value of FoM occurring when $\theta=\pi/2$. The maximum value for the FoM is also 35% greater when initialising qubit B with $\phi=\pi$ with the system demonstrating a better performance when introducing initial coherences.
Figure 8: First row: Figure of merit of the total system in units of $\epsilon_{2B}$ for $0\ \leq\ \theta\ \leq\ \pi$ (x-axis) and $0\ \leq$ $\phi$ $\ \leq\ 2\pi$ (y-axis) for two-qubit systems where columns from left to right are case 1, 2, and 3 respectively. Parameters: $\epsilon_{1j}=0\;\epsilon_{2B},\ \epsilon_{2j}=1\;\epsilon_{2B}$ for $j=A,B$, and $C$. $k_{B}T_{A}=4\;\epsilon_{2B}$ and $k_{B}T_{C}=0.4\;\epsilon_{2B}$. Brighter shades correspond to a greater value of the figure of merit. Second row: Same parameters as the first row but for three-qubit systems.
@@ V Comparison with overall thermal initialization
In previous sections, we have seen that initial coherences tend to favor higher performances. Here we compare the systems analysed so far with systems in which qubit B is initialized in a thermal state. Our results show clear advantage of having qubit B initialized in a pure state.
In Fig. 9 dark and light blue diamonds correspond to the original systems, i.e. with B initialised in a pure state, for three and two qubit systems, respectively, while red and orange diamonds map the systems with B initially thermal (Bthermal systems), for three and two qubit systems, respectively. The figure shows the results for the optimal initial configurations producing ergotropy (first panel), maximum ergotropy generation (second panel), maximum ergotropy ratio (third panel), and maximum FoM (fourth panel) for case 1, case 2, case 3, as labelled. Qubits $A$ and $C$ are always prepared in the thermal state eq. 1 with fixed temperatures $k_{B}T_{A}=4\;\epsilon_{2B}$ and $k_{B}T_{C}=0.4\;\epsilon_{2B}$ (only $k_{B}T_{A}$ applies when dealing with a two-qubit system). We note that each point in a panel may correspond to different initial condition, depending on which initial value of $\theta$ and $\Phi$ will give the overall maximum of that particular quantity over 30 iterations.
@@
The systematic advantage we observe in the original systems (blue and light blue diamonds) comes from the quantum coherences in the initial state of qubit B, as can be seen, e.g., from Fig. 7 and Fig. 8. This quantum advantage translates into a maximum ergotropy (first panel) about an order of magnitude higher for case 2, and two to five times higher for case 1 and 3 .
@@
Figure 9: Comparison of performance measures: Maximum ergotropy $W_{max}$ (first panel), maximum variation in ergotropy $\Delta W_{max}$ (second panel), maximum ergotropy ratio $W_{ratio}$ (third panel) and maximum figure of merit FoM (fourth panel) of the total system for case 1 (C1), case 2 (C2), case 3 (C3), as labelled. Three-qubit system (blue diamonds), two-qubit system (light-blue), Three-qubit system with Bthermal (red) , two-qubit system with Bthermal (orange). Parameters: $k_{B}T_{A}=4\;\epsilon_{2B}$ and $k_{B}T_{C}=0.4\;\epsilon_{2B}$. Maxima calculated over two cycles with $N=15$, $0\leq\phi\leq 2\pi$ and $0\leq\theta\leq\pi$.
@@ VI Experimental feasibility
The protocol requires several ingredients, each of them will be discussed separately: it is not challenging for any existing trapped ion quantum processor to operate a small set of 2 or 3 qubits. The initialization of qubit B into a superposition state is also a simple task for any quantum processor. The inizialization to Gibbs states of qubits A and C, however, is a non-standard operation, but it has been realized in some trapped ion quantum setups [4, 31, 37]. The step-wise execution of the gate, namely the Nth root two-qubit gate, requires accurate calibration of the qubit diving, possibly a challenge when N is large. However, a step-wise execution appears feasible in trapped ion quantum processing, since recently it has been realized but for a single qubit operation [31] with N up to 10, and it is also well established for trotterized simulations like for example in the trapped ion quantum simulation of a high energy system [38]. Finally, for the characterization of the quantum battery output, all methods for state tomography are readily available in trapped ions but also in any of the other platforms for quantum processors.
@@ VII Discussion
A benefit of this current investigation is that NRCG and the two-three qubit circuits we propose could be realised experimentally on a number of different platforms by using computational gates; this is an important next step in the development of quantum batteries as noted by Campaioli et al [20]. When looking at other investigations that inspired this work the focus is generally on maximum work extraction which we first looked at in section. IV.1. The importance of initial coherences for the ergotropy was noted Refs [39, 40]. While we find that the ergotropy is positively influenced by initial coherence, the other measure of performance we consider are even more affected, as summarized below. The power of charging and by extension the efficiency [19, 41, 42, 43] are also used in quantum battery research. The power is investigated in the appendix B.
Our investigation compares a set of 3 different protocols with a two or three identical qubits as working fluid. One of the qubit is initialized in a pure state. To establish which protocol and system is best as battery under which circumstances, has led us to consider a set of measures of performance. Hence, we focus not only on how much work or power can be extracted, but also on which fraction of the system internal energy is extractable and characterise this by the ergotropy ratio (eq.10), used also in [44]. We find that regions with better performance depend on the initialisation of the system just as much as the method used for charging, generally favouring initial qubit B states where $\theta>\pi/2$ and hence with some degree of population inversion.
The difference between initial and final ergotropies $\Delta W$ is introduced as a performance indicator. When looking at optimising initial conditions for battery charging this is arguably a metric more important than just the total ergotropy, especially when initialising the system in a state other than the ground state. With this we can discriminate clear winners among the cases implemented, with case 2 under-performing against cases 1 and 3. This indicator also clarifies charging and discharging regimes with positive regions corresponding to charging and negative to discharging. We find that initial coherence in qubit B is definitely beneficial to charging, improving the maximum ergotropy difference by 26.5% (compare results for $\phi=0$ and $\phi=\pi$). Results for the power in appendix B confirm the importance of initial coherences for enhanced performances.
@@
In the last part of the paper we compare performances for all protocols with corresponding systems with qubit B initialized in a thermal state. The main takeaway from these comparisons is that introducing a pure-state initial component is advantageous across the board, and for all measures of performance, see Fig. 9. Initial coherences lead to overall maxima for some measures of performance.
Further investigations could expand this work by utilizing alternative Nth root computational gates and different quantum thermal machines. We found that, depending on protocol, the best performer could be either the two or the three qubit system: future work could explore the scaling (and its consistency) for increasing number of qubits.
VIII Conclusion
In this paper we have shown that circuits containing Nth-root CNOT gates could be advantageously used as quantum batteries. We evaluate the circuits’ performance using five different measures – ergotropy, power, ergotropy difference and ratio, and a figure of merit – and considering two and three qubit circuits. Our results show consistently that having one of the qubits initialized in a pure state is highly advantageous with respect to using a thermal state with the same energy: high performances are strongly and positively influenced by initial quantum coherences. For any given initial condition, and at difference with standard CNOT gates, iterating circuits with Nth-root CNOT gates allows access to a fine distribution of values of ergotropy: this flexibility could be advantageous for tailoring the same battery to different working needs. The systems and protocol we propose are experimentally feasible with current technology.

### v2

@@ IV.2 Comparison Between Circuits with Full CNOT and NRCGs
NCRGs can be interpreted as a stepped application of a full CNOT gate as explained more thoroughly in Section II.2. We then wish to compare whether applying fractions of CNOTs instead of a full CNOT is advantageous in the present case. We compare the ergotropy over two cycles in which iterations have either of one CNOT or N = 15 steps of an NRCG, which is magenta. For each panel we consider $\theta$ such that the ergotropy of the system considered in the panel reaches its maximum when N = 15. The top left and top center panels in Figure 3, showing case 1 and 2, for the two-qubit systems operate with only a single gate per iteration. These particular systems demonstrate well how the NRCGs operate as a slower application of the full CNOT gate, with the systems’ ergotropy reaching the same values for both types of gates after one cycle. Circuits with the full CNOT gate allow us to reach only two specific values of ergotropy. However, NRCGs allow access to these and to all intermediate values. Such fine-tuning (increase and decrease) of ergotropy with the system’s evolution could prove advantageous when a particular value of ergotropy or the corresponding intermediate state are desired that are in-between the initial and final ergotropy and states generated by a full CNOT gate.
Figure 3: Comparing the ergotropy $W_{max}$ for the NCRG-protocol with $0\ \leq$ N $\ \leq\ 30$ (magenta), and the standard CNOT (blue), in the case of the two-qubit systems (first row) and the three-qubit system (second row), where columns from left to right are cases 1, 2, and 3, respectively. The Y-axis scale differs for each panel. The chosen value of $\theta$ is indicated in each panel. Parameters: $\Phi=\pi$, $\epsilon_{1j}=0\epsilon_{2B},\ \epsilon_{2j}=1\epsilon_{2B}$ for $j=A,B$, and $C$; $k_{B}T_{A}=4\epsilon_{2B}$ and $k_{B}T_{C}=0.4\epsilon_{2B}$. Magenta line: N = 15; blue line: N = 1.
@@
With the exception of cases 1 and 2 for two-qubit systems, NRCG-based cycles are not equivalent to cycles from corresponding circuits with CNOTs, even when considering the end results, see Figure 3. Apart from the easy reach of intermediate values of ergotropy, we can see two other ways in which the use of NRCGs could be advantageous: (i) at points where the ergotropy of the total system has a larger value than that of the standard CNOT gate within the same cycle; or (ii) when the maximum ergotropy reached using NRCGs is higher that the one of the corresponding CNOT gate circuits. The first possibility can be observed in all panels of Figure 3; the second can be seen in the three-qubit cases 1 and 2.
@@
Similar results are found for the case of qubit B, initialised in $\Phi=0$. We note that a larger value of the ergotropy may be reached after the two cycles considered in Figure 3.
@@ IV.3 Ergotropy Variation for Two- and Three-Qubit Systems
Our systems do not start from a passive state, so the ergotropy $W_{max}$ at the end of the process may be influenced by the initial one. This is not a simple relation, as the systems studied undergo ‘discharging’ and ‘charging’ processes as the number of iterations increase, as discussed below.
Figure 4: Ergotropy variation $\Delta W_{max}$ for $\phi=0$. Otherwise, we have the same parameters as in Figure 2.
Figure 5: Ergotropy variation $\Delta W_{max}$ for $\phi=\pi$, with the same parameters as in Figure 2.
In addition, when considering $W_{max}$ (Figure 2), a direct comparison between the two- and three-qubit systems is not easy to perform: when all common parameters are the same, the three-qubit system inherently starts from more energy and ergotropy than the two-qubit system by merit of having more qubits. For the dynamics considered, this almost always translates into acquiring more ergotropy by the end of the process (compare the first and second row in Figure 2). From this, we can argue that looking at the maximum ergotropy may not be a fair comparison between systems with different quantities of qubits. For these reasons, we introduce the ergotropy variation ($\Delta W_{max}$) which is defined by the following equation:
@@
where $W_{max}^{(f)}$ is the ergotropy of the final state and $W_{max}^{(0)}$ is the ergotropy of the initial state of the total system. Equation (9) allows the identification of two regimes with respect to the initial state, charging or discharging. A charging regime is characterised by a positive value of ergotropy variation and a discharging regime is characterised by a negative value of ergotropy variation. $\Delta W_{max}$ allows for a clearer comparison across different systems. For example, an equal or greater positive change in the two-qubit system compared to the corresponding three-qubit system may suggest that an easier-to-implement two-qubit circuit may be more suited to accrue energy as a battery.
[…]
@@ IV.4 Ergotropy Ratio for Two- and Three-Qubit Circuits
Here, we introduce the ergotropy ratio which is defined as,
@@
where $W_{max}$ is the ergotropy (Equation (7)) and $Tr[\rho^{(f)}H]$ is the internal energy of the system in its final state. The ergotropy ratio takes a value between 0 and 1, where a value closer to 1 signifies a larger quantity of extractable energy. The ergotropy ratio is a good measure of how efficiently a battery is performing [39]. With this in mind, here, $W_{ratio}$ is introduced to look for optimal initial conditions for an efficient quantum battery.
Figure 6 shows the ergotropy ratio for the total system for the same parameters as those in Figure 2, in particular, qubit B is initialised with $\Phi=\pi$. For all cases, we obtain a greater value of ergotropy ratio when qubit B is initialised with a large value of $\theta$, but not necessarily $\theta=\pi$. The maximum values for $W_{ratio}$ across all panels in Figure 6 are similar, suggesting no particular advantage with using a three- instead of a two-qubit system. When initialising qubit B with $\Phi=\pi$ (see Figure 6), we observe a marked increased in areas of high $W_{ratio}$ for values of $\theta$ closer to $\pi/2$ for both two- (case 3) and three- (cases 1 and 3) qubit systems, supporting the advantage from a larger initial coherence. For comparison, results with qubit B initialised at $\Phi=0$ are shown in Figure 14, Appendix C.
Figure 6: Ergotropy ratio $W_{ratio}$ for $\Phi=\pi$, with the same parameters as in Figure 2.
While $W_{ratio}$ is a good indicator for which initial conditions lead to a larger percentage of total energy that can be extracted as work, it does not show if that large percentage also corresponds to a region of high charging: indeed, $W_{ratio}$ and $\Delta W_{max}$ have quite different behaviours (e.g., compare Figures 4 and 14). Ideally, one would like to optimize both quantities at the same time to quantitatively assess how well the initial conditions, system, and circuit considered perform as a battery. In the next section, we will introduce a related figure of merit.
@@ IV.5 Figure of Merit for Optimal Battery, Two- and Three-Qubit Systems
The figure of merit we propose to find the optimal circuit, number of iterations, and initial conditions for charging a quantum battery has the form
@@
This combination of quantities shows where the best balance of energy that can be extracted (ergotropy ratio) and the ergotropy increase from the initial condition (ergotropy difference) lie, giving a clearer indication of the optimal system.
In Figure 7, we plot $FoM$ for the same parameter scans of Figures 5 and 6. In particular, qubit B is initialised with $\Phi=\pi$. The results corresponding to $\Phi=0$ are plotted in Appendix C, Figure 15. The topography of these graphs tends to mirror that of the variation in ergotropy, which is in general more variable with iterations than that of the ergotropy ratio. Interestingly, when $\Phi=0$ (Figure 15), we see the highest FoM value regions when $\theta\approx 0$ and $\theta\approx\pi$, suggesting that the system performs better as a battery when qubit B is initialised closer to a state with no coherences ($\Phi=0$, with $\theta=0$ or $\pi$). When comparing this to Figure 7 ($\Phi=\pi$), the opposite is true, with the greatest value of FoM occurring when $\theta=\pi/2$. The maximum value for the FoM is also 35% greater when initialising qubit B with $\phi=\pi$ with the system demonstrating a better performance when introducing initial coherences.
Figure 7: Figure of merit, FoM, with $\Phi=\pi$, with the same parameters as Figure 2.
@@
Figure 8: First row: Figure of merit of the total system in units of $\epsilon_{2B}$ for $0\ \leq\ \theta\ \leq\ \pi$ (x-axis) and $0\ \leq$ $\phi$ $\ \leq\ 2\pi$ (y-axis) for two-qubit systems where columns from left to right are case 1, 2, and 3, respectively. Parameters: $\epsilon_{1j}=0\epsilon_{2B},\ \epsilon_{2j}=1\epsilon_{2B}$ for $j=A,B$, and $C$. $k_{B}T_{A}=4\epsilon_{2B}$ and $k_{B}T_{C}=0.4\epsilon_{2B}$. Brighter shades correspond to a greater value of the figure of merit. Second row: The same parameters as the first row but for three-qubit systems.
@@ V Comparison with Overall Thermal Initialization
In previous sections, we have seen that initial coherences tend to favor higher performances. Here, we compare the systems analysed so far with systems in which qubit B is initialised in a thermal state. Our results show a clear advantage of having qubit B initialised in a pure state.
In Figure 9, dark and light blue diamonds correspond to the original systems, i.e.,  with B initialised in a pure state, for three- and two-qubit systems, respectively, while red and orange diamonds map the systems with B initially being thermal (Bthermal systems) for three- and two-qubit systems, respectively. The figure shows the results for the optimal initial configurations producing ergotropy (first panel), maximum ergotropy generation (second panel), the maximum ergotropy ratio (third panel), and maximum FoM (fourth panel) for case 1, case 2, and case 3, as labelled. Qubits $A$ and $C$ are always prepared in the thermal state Equation (1) with fixed temperatures $k_{B}T_{A}=4\epsilon_{2B}$ and $k_{B}T_{C}=0.4\epsilon_{2B}$ (only $k_{B}T_{A}$ applies when dealing with a two-qubit system). We note that each point in a panel may correspond to a different initial condition, depending on which initial value of $\theta$ and $\Phi$ will give the overall maximum of that particular quantity over 30 iterations.
Figure 9: Comparison of performance measures: maximum ergotropy $W_{max}$ (first panel), maximum variation in ergotropy $\Delta W_{max}$ (second panel), maximum ergotropy ratio $W_{ratio}$ (third panel), and maximum figure of merit FoM (fourth panel) of the total system for case 1 (C1), case 2 (C2), and case 3 (C3), as labelled. Three-qubit system (blue diamonds), two-qubit system (light-blue), Three-qubit system with Bthermal (red) , two-qubit system with Bthermal (orange). Parameters: $k_{B}T_{A}=4\epsilon_{2B}$ and $k_{B}T_{C}=0.4\epsilon_{2B}$. Maxima calculated over two cycles with $N=15$, $0\leq\phi\leq 2\pi$ and $0\leq\theta\leq\pi$.
@@
The systematic advantage we observe in the original systems (blue and light blue diamonds) comes from the quantum coherences in the initial state of qubit B, as can be seen, e.g., from Figures 7 and 8. This quantum advantage translates into a maximum ergotropy (first panel) about an order of magnitude higher for case 2, and two to five times higher for cases 1 and 3 .
@@ VI Experimental Feasibility
The protocol requires several ingredients, each of which will be discussed separately: it is not challenging for any existing trapped ion quantum processor to operate a small set of 2 or 3 qubits. The initialization of qubit B into a superposition state is also a simple task for any quantum processor. The inizialization to Gibbs states of qubits A and C, however, is a non-standard operation, but it has been realized in some trapped ion quantum setups [4, 30, 40]. The step-wise execution of the gate, namely the Nth root two-qubit gate, requires the accurate calibration of the qubit diving, which is possibly a challenge when N is large. However, a step-wise execution appears to be feasible in trapped ion quantum processing, since recently it has been realized but for a single qubit operation [30] with N up to 10, and it is also well established for trotterized simulations like, for example, in the trapped ion quantum simulation of a high-energy system [41]. Finally, for the characterization of the quantum battery output, all methods for state tomography are readily available in trapped ions but also in any of the other platforms for quantum processors.
@@ VII Discussion
A benefit of this current investigation is that NRCGs and the two–three qubit circuits we propose could be realised experimentally on a number of different platforms by using computational gates; this is an important next step in the development of quantum batteries, as noted by Campaioli et al. [20]. When looking at other investigations that inspired this work, the focus is generally on maximum work extraction, which we first looked at in Section IV.1. The importance of initial coherences for the ergotropy was noted Refs. [42, 43]. While we find that the ergotropy is positively influenced by initial coherence, the other measures of performance we consider are even more affected, as summarized below. The power of charging, and by extension the efficiency [19, 44, 45, 13], are also used in quantum battery research. The power is investigated in Appendix B.
Our investigation compares a set of three different protocols with two or three identical qubits as working fluid. One of the qubits is initialised in a pure state. The idea to establish which protocol and system is best as a battery under which circumstances has led us to consider a set of measures of performance. Hence, we focus not only on how much work or power can be extracted, but also on which fraction of the system’s internal energy is extractable and characterised this by the ergotropy ratio (Equation (10)), used also in [39]. We find that regions with better performance depend on the initialisation of the system just as much as the method used for charging, generally favouring initial qubit B states where $\theta>\pi/2$ and hence with some degree of population inversion. However, comparison with results from corresponding initial states where the off-diagonal elements of qubit B are set to 00 demonstrates that initial coherences significantly enhance the ergotropy ratio for $\theta<\pi/2$.
The difference between initial and final ergotropies $\Delta W$ is introduced as a performance indicator. When looking at optimising initial conditions for battery charging, this is arguably a metric that is more important than just the total ergotropy, especially when initialising the system in a state other than the ground state. With this, we can discriminate clear winners among the cases implemented, with case 2 under-performing against cases 1 and 3. This indicator also clarifies charging and discharging regimes, with positive regions corresponding to charging and negative to discharging. We find that initial coherence in qubit B is definitely beneficial to charging, improving the maximum ergotropy difference by 26.5% (compare results for $\phi=0$ and $\phi=\pi$). Results for the power in Appendix B confirm the importance of initial coherences for enhanced performances.
@@
In the last part of this paper, we compare performances for all protocols with corresponding systems with qubit B initialised in a thermal state. The main takeaway from these comparisons is that introducing a pure-state initial component is advantageous across the board, and for all measures of performance, see Figure 9. Initial coherences lead to overall maxima for some measures of performance.
Further investigations could expand this work by exploring the effect of intermediate measurements on the protocols’ efficiency [46] and utilizing alternative Nth root computational gates and different quantum thermal machines. We found that, depending on the protocol, the best performer could be either the two- or the three-qubit system: future research could explore the scaling (and its consistency) for increasing the number of qubits. Another avenue of future investigation could be the use of a collective bath [47] and of global collective charging [48] to verify if, in our case, they would improve performances.
A challenge that this system and others could face experimentally is the loss of coherences, whether from dissipation or through direct measurement [46]. How ergotropy and related measures of performance would be affected would depend on the dissipation mechanisms specific to the hardware. However, as hardware for quantum computers has already been optimized to perform a very large number of gates within the relevant decoherence times, quantum batteries based on quantum circuits should also share this advantage.
VIII Conclusions
In this paper we have shown that circuits containing Nth-root CNOT gates could be advantageously used as quantum batteries. We evaluate the circuits’ performance using five different measures—ergotropy, power, ergotropy difference and ratio, and a figure of merit—and considering two- and three-qubit circuits. Our results show consistently that having one of the qubits initialised in a pure state is highly advantageous with respect to using a thermal state with the same energy: high performances are strongly and positively influenced by initial quantum coherences. For any given initial condition, and at difference with standard CNOT gates, iterating circuits with Nth-root CNOT gates allows access to a fine distribution of values of ergotropy: this flexibility could be advantageous for tailoring the same battery to different working needs. The systems and protocol we propose are experimentally feasible with the current technology.